# Efficient Generative Model Training via Embedded Representation Warmup

### v1

@@ Abstract
Diffusion models excel at generating high-dimensional data but fall short in training efficiency and representation quality compared to self-supervised methods. We identify a key bottleneck: the underutilization of high-quality, semantically rich representations during training notably slows down convergence. Our systematic analysis reveals a critical representation processing region—primarily in the early layers—where semantic and structural pattern learning takes place before generation can occur. To address this, we propose Embedded Representation Warmup (ERW), a plug-and-play framework where in the first stage we get the ERW module serves as a warmup that initializes the early layers of the diffusion model with high-quality, pretrained representations. This warmup minimizes the burden of learning representations from scratch, thereby accelerating convergence and boosting performance. Our theoretical analysis demonstrates that ERW’s efficacy depends on its precise integration into specific neural network layers—termed the representation processing region—where the model primarily processes and transforms feature representations for later generation. We further establish that ERW not only accelerates training convergence but also enhances representation quality: empirically, our method achieves a 40$\times$ acceleration in training speed compared to REPA [54], the current state-of-the-art methods. Code is available at https://github.com/LINs-lab/ERW.
Figure 1: Embedded Representation Warmup (ERW). Throughout the training process, we demonstrate that incorporating representations at the early stages is highly beneficial. To this end, we propose a representation warmup stage that employs a representation alignment loss to integrate representations from models such as Dinov2  [38] into the ERW. This initialized representation region is subsequently embedded into the diffusion model pipeline, providing a strong starting point for training. Our ERW method thus enhances both efficiency and effectiveness, leading to faster convergence and superior performance compared to the REPA [54] method, thereby establishing a new state-of-the-art.
@@ 1 Introduction
Deep generative models, particularly diffusion models [16, 49], have made remarkable strides in producing high-fidelity, realistic images. These models have demonstrated impressive performance in tasks such as unconditional image generation [8] and text-to-image synthesis [43, 45], largely due to their capacity to model complex data distributions through iterative denoising processes. By refining noisy inputs over multiple stages, diffusion models effectively capture intricate features that yield visually coherent outputs.
Despite these achievements, diffusion models often lag behind self-supervised learning methods in both training efficiency and the quality of learned representations. This gap is notable because efficient training and robust representation learning are pivotal for scaling generative models to real-world applications. While diffusion models implicitly learn semantic features during the denoising process [53, 52], these representations typically lack the depth and versatility seen in self-supervised approaches [3, 38]. Recent studies [29, 54] have investigated techniques to enhance internal representations in diffusion models, aiming to improve both the expressiveness of the generated data and the training efficiency. Moreover, Kadkhodaie et al. [20] highlight a critical bottleneck: the tension between memorizing semantic information and generalizing to realistic data distributions.
These observations lead to a pivotal question:
$\mathbb{Q}$: Can we first learn high-quality external representations from a well-established representation model and then integrate their representation into diffusion models, rather than training both representation and generation from scratch?
@@
Self-supervised learning approaches, such as contrastive learning [4] and masked autoencoders [14], as well as more recent methods like Dinov2 [38], have shown strong capabilities in extracting rich semantic representations from unlabeled data. However, adapting these external representations to diffusion models poses inherent challenges. Primarily, diffusion models operate on progressively denoised inputs, whereas self-supervised encoders are trained on clean data. In addition, the architectures of self-supervised encoders may differ significantly from typical diffusion model backbones, further complicating direct integration.
Our approach. We observe that diffusion model training can be effectively conceptualized as a two-phase process: first learning semantic representations, and then leveraging those representations for generative purposes. Motivated by this viewpoint, we propose Embedded Representation Warmup (ERW), which explicitly separates training into two phases. In the warmup phase, we initialize the early-layer parameters of the diffusion model with weights from a pretrained self-supervised backbone (e.g., Dinov2). Since early layers predominantly capture semantic features, this partial initialization effectively injects strong semantic priors into the diffusion model. Our ablation studies confirm that restricting representation learning to early layers, rather than extending it to the entire architecture, leads to both better performance and higher training efficiency.
Furthermore, we decompose the diffusion model’s workflow into two distinct circuits: ① the Latent-to-Representation (L2R) circuit, mapping raw inputs into a semantic embedding space, and ② the Representation-to-Generation (R2G) circuit, synthesizing final outputs from these embeddings. Building on this insight, ERW uniformly initializes the model’s early layers to embed robust semantic features, forming a solid foundation for subsequent generative training.
Validation. Extensive experiments confirm that providing high-quality representations from the outset via ERW substantially accelerates convergence. We demonstrate a significant training speedup (up to $\mathbf{40\times}$ compared to REPA) while maintaining state-of-the-art generative quality. Notably, ERW enables SiT-XL to reach an FID of 6.4 on ImageNet-1k without classifier-free guidance in under 4 hours on 8$\times$×H800 GPUs. Since the warmup phase accounts for only a fraction of the overall training cost, reusing warmup-optimized checkpoints further enhances training efficiency across multiple tasks.
Contributions:
@@
    We identify and demonstrate that diffusion model training can be effectively decomposed into two distinct phases: a representation learning phase and a generation phase, with the representation processing mainly occurring in early layers (see  Section 3).
@@
    We introduce Embedded Representation Warmup, a two-phase training framework that first initializes the representation layers using pretrained self-supervised models before proceeding to the main generation-focused training (see  Section 4).
@@
    We empirically validate our approach on and demonstrating significant improvements in training efficiency and generation quality (see  Section 5).
@@ 2 Related Work
Our work builds upon and extends several areas of research in diffusion models, representation learning, and the integration of external representations for generative modeling.
Integrating pretrained representations into diffusion models.
Recent work exploits pretrained encoders to enhance diffusion-based generation. Würstchen [41] adopts a two-stage design where an initial network learns a semantic map, guiding a second diffusion model for high-resolution output. RCG [32] further underscores the utility of compact latent codes, generated by a first diffusion model and consumed by a second, boosting generative performance. REPA [54] applies regularization throughout the diffusion training process to align model representations with pretrained visual encoders. In contrast, our approach builds on these ideas by decoupling representation learning from the denoising process. We explicitly separates the training into two phases first learning representation then learning generation. This design not only alleviates the optimization challenges associated with joint learning but also leads to a significant acceleration in training convergence.
Leveraging self-supervised learning for generative modeling.
Self-supervised encoders pretrained via contrastive learning [4] or masked reconstruction [14] have made significant inroads in representation quality. Recent developments integrate such encoders into generative pipelines, for instance using them as discriminators in GAN-based frameworks [47, 26] or to guide diffusion distillation [30].
Enhancing representation learning within diffusion models.
Numerous studies highlight the discriminative potential of diffusion representations [53, 52], revealing that intermediate layers encode semantic features. Efforts like Repfusion [53] and DreamTeacher [30] distill these features to smaller networks for downstream tasks. Inspired by these findings, we identify a distinct representation processing region where crucial representation learning occurs before generation can take place. By initializing this region with pretrained weights during a dedicated warmup phase, we circumvent the costly need to learn complex features from scratch in the main training phase, leading to faster convergence and improved fidelity.
Towards unified and efficient diffusion training.
Beyond two-stage pipelines [41, 32], alternative methods fuse self-supervision with the diffusion objective, such as MaskDiT [55] or MAGE [31], yet often require careful objective balancing. Our plug-and-play method instead introduces a clear two-phase approach: a representation-focused warmup phase that initializes early diffusion layers with mature pretrained weights, followed by a standard diffusion training phase that benefits from this strong representational foundation.
@@ 3.1 Preliminaries
While classic diffusion models such as DDPM [16] adopt a discrete-time denoising process, flow-based methods [33, 1, 48] explore diffusion in a continuous-time setting. In particular, Scalable Interpolant Transformers (SiT) [10, 33, 34] offer a unifying framework for training diffusion models on a continuous-time stochastic interpolant. Below, we describe how SiT can be leveraged to learn powerful latent diffusion models.
@@
Consider a data sample $\mathbf{x}\sim p(\mathbf{x})$ (e.g., an image) and let the encoder $\mathcal{H}_{\bm{\theta}}(\mathbf{x})$ map it to its latent representation denoted as $\mathbf{z}_{\ast}\in\mathcal{Z}$. Given standard Gaussian noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, SiT defines a forward process in the latent space, parameterized by continuous time $t\in[0,1]$:
$\mathbf{z}_{t}=\alpha_{t}\,\mathbf{z}_{\ast}\,+\,\sigma_{t}\,\bm{\epsilon}\,,$ (1)
@@
This construction implies that $\mathbf{z}_{0}=\mathbf{z}_{\ast}$ (the clean latent representation) and $\mathbf{z}_{1}=\bm{\epsilon}$ (pure noise). Under mild conditions [1], the sequence $\{\mathbf{z}_{t}\}$ forms a stochastic interpolant that smoothly transitions between data and noise in the latent space.
@@
$\dot{\mathbf{z}}_{t}=\dot{\alpha}_{t}\,\mathbf{z}_{\ast}\,+\,\dot{\sigma}_{t}\,\bm{\epsilon}\,.$ (3)
@@
$\small\textstyle\mathcal{L}_{\mathrm{velocity}}(\bm{\theta})\;=\;{\mathbb{E}}_{\mathbf{z}_{\ast},\bm{\epsilon},t}\left[\left\lVert\mathbf{v}_{\bm{\theta}}(\mathbf{z}_{t},t)-\Bigl{(}\dot{\alpha}_{t}\,\mathbf{z}_{\ast}+\dot{\sigma}_{t}\,\bm{\epsilon}\Bigr{)}\right\rVert^{2}\right]\,.$ (5)
Successfully learning $\mathbf{v}_{\bm{\theta}}(\mathbf{z},t)$ enables one to integrate the reverse-time ordinary differential equation (ODE) [49], thereby mapping noise samples back to coherent latent representations.
@@ 3.2 Three-Stage Diffusion Circuit
Recent studies indicate that diffusion models jointly perform both representation learning and generative decoding during the denoising procedure [54, 52]. Notably, every layer in the network contributes to feature extraction and generative tasks to varying degrees. To make this dual functionality clearer, we propose decomposing the diffusion process into three distinct stages: Pixel-to-Latent (P2L), Latent-to-Representation (L2R), and Representation-to-Generation (R2G), as illustrated in Figure 2. Formally, we posit that the diffusion sampling procedure can be written as:
@@
Here, the Variational Autoencoder (VAE) encoder $\mathcal{H}_{\bm{\theta}}$ compresses high-dimensional pixel data $\mathbf{x}$ into a lower-dimensional latent $\mathbf{z}$. The diffusion model components $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$ and $\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}$ share parameters in diffusion model backbone but serve distinct functions: $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$ extracts semantic representations from noisy latent codes (L2R), while $\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}$ transforms these representations into refined latent outputs (R2G). During inference, a VAE decoder $\mathcal{D}_{\bm{\theta}}$ would convert these refined latents back to pixel space. This decomposition into distinct processing stages provides a more transparent view of how latent diffusion models function and support the idea from latent space compression in Saharia et al. [45] and the diffusion model training via represention alignment in REPA Yu et al. [54].
@@
The diffusion model’s objective is to learn the score function $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})$, which represents the gradient of the log probability density of the data distribution. Our key theoretical insight is that this score function can be orthogonally decomposed into representation and generation components (see details in Appendix A):
@@
This decomposition permits us to partition the overall diffusion objective into two nearly orthogonal terms, focusing respectively on (i) extracting meaningful features from corrupted inputs and (ii) translating those features into coherent outputs. Such a structure naturally aligns with the three-stage diffusion circuit we will soon discuss, clarifying the distinct roles played by different network regions.
@@
Before performing the denoising process in the high-dimensional pixel domain—where noise may obscure semantic cues—many contemporary methods [45, 16, 8] compress images into a more tractable latent space:
@@
Figure 3: Selected Samples on ImageNet $256\times 256$. Images generated by the SiT-XL/2 + REPA + ERW model using classifier-free guidance (CFG) with a scale of $w=2.2$ under 40 epochs.
@@
$\mathbf{r}_{t}\;=\;\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}(\mathbf{z}_{t})\,.$ (8)
This step corresponds to estimating $\nabla\log p(\mathbf{r}_{t})$ within the velocity loss decomposition (where we separate representation learning from generation, as detailed in  Appendix A), sometimes termed the Representation Score [49]. Intuitively, the model should discern salient patterns (e.g., object shapes, style characteristics, or conditioning signals) before denoising. Under the sufficient statistic assumption (see Assumption 1), the representation $\mathbf{r}_{t}$ effectively captures the essential information from the latent $\mathbf{z}_{t}$, allowing us to formulate the representation learning process as an independent objective:
$\mathcal{L}_{\mathrm{L2R}}\;=\;\mathbb{E}_{t,\mathbf{z}_{t}}\bigl{[}\|\,\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}(\mathbf{z}_{t})\!-\!\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})\,\|^{2}\bigr{]}\,,$ (9)
@@
In the final phase of each reverse diffusion update in (3), known as the R2G stage, the extracted semantic representation is transformed into an updated latent with reduced noise:
@@
$\mathbf{r}_{t}=\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}(\mathbf{z}_{t})\quad\longrightarrow\quad\mathbf{z}_{t-\Delta t}=\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}(\mathbf{r}_{t})$ (12)
Hence, every time step naturally splits into (i) L2R for refining the representation and (ii) R2G for synthesizing an updated latent. This loop neatly implements the principle of “first representation, then generation”. Empirically, prior work [54, 52] confirms that early layers of the diffusion model predominantly focus on representation extraction, whereas later layers emphasize generative refinement. Consequently, the staged design mirrors the reverse-time diffusion trajectory, concluding in a final latent $\mathbf{z}_{0}$ that is decoded via $\mathcal{D}_{\bm{\theta}}$ to yield the synthesized output $\mathbf{x}_{0}$.

### v2

@@ Abstract
Generative models face a fundamental challenge: they must simultaneously learn high-level semantic concepts (what to generate) and low-level synthesis details (how to generate it). Conventional end-to-end training entangles these distinct, and often conflicting objectives, leading to a complex and inefficient optimization process. We argue that explicitly decoupling these tasks is key to unlocking more effective and efficient generative modeling. To this end, we propose Embedded Representation Warmup (ERW), a principled two-phase training framework. The first phase is dedicated to building a robust semantic foundation by aligning the early layers of a diffusion model with a powerful pretrained encoder. This provides a strong representational prior, allowing the second phase—generative full training with alignment loss to refine the representation—to focus its resources on high-fidelity synthesis. By separating the learning of "what" from "how," ERW fundamentally simplifies the optimization landscape. Our analysis confirms that this efficacy stems from functionally specializing the model’s early layers for representation. Empirically, our framework achieves up to a 11.5$\times$ speedup in training iterations to reach FID=1.41 compared to single-phase methods like REPA (Yu et al., 2024). Code is available at https://github.com/LINs-lab/ERW.
Figure 1: A Staged Approach: First Build Semantics, Then Synthesize. Our framework operationalizes the decoupling of semantic understanding from generative synthesis. In Phase 1 (Semantic Foundation), we exclusively train the model’s early layers to align with a pretrained encoder (e.g., DINOv2 (Oquab et al., 2023)), establishing a robust understanding of what to generate. In Phase 2 (Guided Synthesis), the full model is trained. The plot empirically demonstrates the power of this decoupling: ERW converges dramatically faster and achieves superior performance compared to single-phase training like REPA (Yu et al., 2024), which entangles both learning tasks.
@@ 1 Introduction
Deep generative models, particularly diffusion models (Ho et al., 2020; Song et al., 2020), have achieved remarkable success in high-fidelity image generation. These models excel at tasks ranging from unconditional image generation (Dhariwal & Nichol, 2021) to text-to-image synthesis (Ramesh et al., 2022; Saharia et al., 2022), demonstrating a profound capacity to model complex data distributions. However, underpinning their impressive capabilities is a fundamental tension, arising from a multitude of entangled learning objectives.
At its core, effective generation requires both semantic understanding—comprehending what constitutes meaningful content—and visual synthesis—translating abstract concepts into precise pixel-level details. Conventional end-to-end training entangles these objectives within a single optimization process, forcing the model to concurrently learn high-level conceptual knowledge and low-level rendering skills. This entanglement creates inherent optimization conflicts, a challenge reminiscent of the classic perception-distortion trade-off (Blau & Michaeli, 2018). Early in training, the model’s attempts to fit pixel-level details may interfere with its ability to capture global semantic structures, an issue exacerbated by the known spectral bias of neural networks towards learning low-frequency components first (Rahaman et al., 2019; Sauer et al., 2021). Consequently, later stages may struggle to refine generation quality due to inadequate representational foundations.
Recent studies have begun to acknowledge this tension. While diffusion models implicitly learn semantic features during denoising (Yang & Wang, 2023; Xiang et al., 2023), these representations often lack the robustness and versatility of dedicated self-supervised approaches (Caron et al., 2021; Oquab et al., 2023). Moreover, Kadkhodaie & Simon (2024) highlight the critical bottleneck between memorizing semantic information and generalizing to realistic distributions. Methods like REPA (Yu et al., 2024) have attempted to address this by aligning diffusion representations with pretrained encoders throughout training, yet they still suffer from the fundamental challenge of joint optimization. These observations lead us to a pivotal question:
$\mathbb{Q}$: Can we fundamentally simplify generative model training by decoupling semantic understanding from visual synthesis, thereby allowing each component to be optimized more effectively?
@@
Self-supervised learning approaches, including contrastive methods (Chen et al., 2020a), masked autoencoders (He et al., 2022), and recent advances like DINOv2 (Oquab et al., 2023), have demonstrated exceptional capabilities in learning rich semantic representations. However, effectively integrating these external representations into diffusion models remains challenging due to fundamental mismatches: diffusion models operate on progressively noisy inputs while self-supervised encoders are trained on clean data, and architectural differences further complicate direct integration.
Our approach. We propose that the key to resolving this challenge lies in explicitly decoupling the learning of semantic understanding from visual synthesis. To this end, we introduce Embedded Representation Warmup (ERW), a principled two-phase framework that operationalizes this decoupling philosophy. Our approach is grounded in the observation that diffusion models naturally exhibit a functional specialization: early layers predominantly handle semantic processing (what we term the Latent-to-Representation or L2R circuit), while later layers focus on generative refinement (the Representation-to-Generation or R2G circuit).
Rather than forcing both circuits to learn simultaneously from scratch, ERW strategically separates their optimization: Phase 1 (Semantic Foundation) establishes a robust semantic foundation by dedicating training exclusively to aligning the L2R circuit with a pretrained self-supervised encoder (e.g., DINOv2). This phase ensures the model is "born in Rome"—equipped with mature semantic understanding from the outset. Phase 2 (Guided Synthesis) then leverages this foundation to focus training resources on the R2G circuit, optimizing visual synthesis under the guidance of a gradually diminishing representational constraint.
Validation. Extensive experiments demonstrate that our decoupling strategy yields substantial benefits. ERW achieves up to an $\mathbf{11.5\times}$ speedup in training iterations compared to single-phase methods like REPA while achieving FID=1.41. The warmup phase requires only a fraction of the total training cost, making our approach highly practical for real-world applications.
Our contributions are threefold:
@@
    We formalize the optimization entanglement in generative models as a conflict between semantic understanding and visual synthesis, and propose a conceptual decomposition of the diffusion model into functionally specialized L2R and R2G circuits.
@@
    We introduce ERW, a principled two-phase training paradigm that operationalizes this decoupling, first building a semantic foundation and then focusing on guided synthesis.
@@
    We demonstrate the effectiveness of our framework through extensive experiments, achieving state-of-the-art results.
@@ 2 Related Work
Our work builds on three research pillars: leveraging pretrained encoders, recent advances in diffusion model acceleration, and enhancing the internal representations of diffusion models through decoupled training strategies.
Leveraging pretrained encoders for guidance.
The idea of leveraging powerful pretrained encoders (Radford et al., 2021; Oquab et al., 2023) to guide generation is well-established, with applications as GAN discriminators (Sauer et al., 2021; Kumari et al., 2022) or for knowledge distillation (Li et al., 2023b). A recent and direct approach is concurrent representation alignment, epitomized by REPA (Yu et al., 2024), which accelerates training by enforcing alignment throughout the entire process. In contrast, our work treats alignment as a foundational warmup, relaxing the constraint during later stages to allow the model to focus fully on synthesis.
Contemporary acceleration strategies and recent advances.
Accelerating diffusion models has emerged as a critical research thrust, as recent years have witnessed significant breakthroughs across multiple fronts (Fuest et al., 2024). Post-training sampling acceleration continues to be actively pursued through knowledge distillation techniques that compress slow teachers into fast students (Salimans & Ho, 2022; Sauer et al., 2023; Shao et al., 2023), and through consistency models enabling one-shot or few-shot generation (Song et al., 2023; Heek et al., 2024). Recent work includes speculative decoding approaches for autoregressive text-to-image generation and training-free acceleration methods. Advanced numerical solvers remain crucial, with improvements to DPM-Solver (Lu et al., 2022) and novel exponential integrators significantly reducing function evaluations. Training acceleration strategies include architectural decoupling in staged pipelines (Karras et al., 2018; Ho et al., 2022; Saharia et al., 2022), curriculum learning on timesteps (Xu et al., 2024), and progressive sparse low-rank adaptation methods. ERW contributes to this rapidly evolving landscape by fundamentally decoupling learning objectives within the training process, separating semantic understanding ("what") from synthesis capability ("how").
Internal vs. injected representations and efficient fine-tuning.
Numerous studies confirm that diffusion models implicitly learn powerful, classifier-like semantic features (Yang & Wang, 2023; Li et al., 2023a; Xiang et al., 2023), a phenomenon some works have deconstructed this phenomenon for self-supervised learning (Chen et al., 2024). An alternative strategy enhances internal representation learning by fusing diffusion objectives with auxiliary self-supervision losses, exemplified by MAGE (Li et al., 2023c) and MaskDiT (Zheng et al., 2024), which draw inspiration from contrastive learning (Chen et al., 2020a; He et al., 2020) and masked autoencoding (He et al., 2022). However, these approaches require careful balancing of competing objectives. ERW sidesteps these complexities by directly injecting mature semantic priors via dedicated warmup, freeing the model to focus purely on high-fidelity synthesis while achieving training efficiency comparable to contemporary acceleration methods.
@@ 3.1 Preliminaries
While classic diffusion models such as DDPM (Ho et al., 2020) adopt a discrete-time denoising process, flow-based methods (Lipman et al., 2022; Albergo et al., 2023; Shi et al., 2024) explore diffusion in a continuous-time setting. In particular, Scalable Interpolant Transformers (SiT) (Ma et al., 2024; Esser et al., 2024; Lipman et al., 2022; Liu et al., 2023) offer a unifying framework for training diffusion models on a continuous-time stochastic interpolant. Below, we describe how SiT can be leveraged to learn powerful latent diffusion models.
@@
Consider a data sample $\mathbf{x}\sim p(\mathbf{x})$ (e.g., an image) and let the encoder $\mathcal{H}_{\bm{\theta}}(\mathbf{x})$ map it to its latent representation denoted as $\mathbf{z}\in\mathcal{Z}$. Given standard Gaussian noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, SiT defines a forward process in the latent space, parameterized by continuous time $t\in[0,1]$:
$\mathbf{z}_{t}=\alpha_{t}\,\mathbf{z}\,+\,\sigma_{t}\,\bm{\epsilon}\,,$ (1)
@@
This construction implies that $\mathbf{z}_{0}=\mathbf{z}$ (the clean latent representation) and $\mathbf{z}_{1}=\bm{\epsilon}$ (pure noise). Under mild conditions (Albergo et al., 2023), the sequence $\{\mathbf{z}_{t}\}$ forms a stochastic interpolant that smoothly transitions between data and noise in the latent space.
@@
$\dot{\mathbf{z}}_{t}=\dot{\alpha}_{t}\,\mathbf{z}\,+\,\dot{\sigma}_{t}\,\bm{\epsilon}\,.$ (3)
@@
$\mathcal{L}_{\mathrm{velocity}}(\bm{\theta})\;=\;{\mathbb{E}}_{\mathbf{z},\bm{\epsilon},t}\left[\left\lVert\mathbf{v}_{\bm{\theta}}(\mathbf{z}_{t},t)-\Bigl{(}\dot{\alpha}_{t}\,\mathbf{z}+\dot{\sigma}_{t}\,\bm{\epsilon}\Bigr{)}\right\rVert^{2}\right]\,.$ (5)
Successfully learning $\mathbf{v}_{\bm{\theta}}(\mathbf{z},t)$ enables one to integrate the reverse-time ordinary differential equation (ODE) (Song et al., 2020), thereby mapping noise samples back to coherent latent representations.
@@ 3.2 Three-Stage Diffusion Circuit
Recent studies indicate that diffusion models jointly perform both representation learning and generative decoding during the denoising procedure (Yu et al., 2024; Xiang et al., 2023). Notably, every layer in the network contributes to feature extraction and generative tasks to varying degrees. To make this dual functionality clearer, we propose decomposing the diffusion process into three distinct stages: Pixel-to-Latent (P2L), Latent-to-Representation (L2R), and Representation-to-Generation (R2G), as illustrated in Figure 2. Formally, we posit that the diffusion sampling procedure can be written as:
@@
Here, the Variational Autoencoder (VAE) encoder $\mathcal{H}_{\bm{\theta}}$ compresses high-dimensional pixel data $\mathbf{x}$ into a lower-dimensional latent $\mathbf{z}$. The diffusion model components $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$ and $\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}$ share parameters in diffusion model backbone but serve distinct functions: $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$ extracts semantic representations from noisy latent codes (L2R), while $\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}$ transforms these representations into refined latent outputs (R2G). During inference, a VAE decoder $\mathcal{D}_{\bm{\theta}}$ would convert these refined latents back to pixel space. This decomposition into distinct processing stages provides a more transparent view of how latent diffusion models function and support the idea from latent space compression in Saharia et al. (2022) and the diffusion model training via representation alignment in REPA Yu et al. (2024).
@@
The diffusion model’s objective is to learn the score function $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})$, which represents the gradient of the log probability density of the data distribution. Our conceptual motivation is that this score function can be orthogonally decomposed into representation and generation components under certain idealizing assumptions (see details in Appendix A):
@@
This decomposition permits us to partition the overall diffusion objective into two nearly orthogonal terms, focusing respectively on (i) extracting meaningful features from corrupted inputs and (ii) translating those features into coherent outputs. While this decomposition is an idealization, as the components are not perfectly orthogonal in a real network and the intermediate representation $\mathbf{r}_{t}$ is not a perfect sufficient statistic, it provides a useful lens for our method. Such a structure naturally aligns with the three-stage diffusion circuit we will soon discuss, clarifying the distinct roles played by different network regions.
@@
Before performing the denoising process in the high-dimensional pixel domain—where noise may obscure semantic cues—many contemporary methods Saharia et al. (2022); Ho et al. (2020); Dhariwal & Nichol (2021) compress images into a more tractable latent space:
@@
Figure 3: Selected Samples on ImageNet $\times 256$ 256. Images generated by the SiT-XL/2 + REPA + ERW model using Classifier-Free Guidance (CFG) with a scale of $w=1.62$ under 350 epochs.
@@
$\mathbf{r}_{t}\;=\;\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}(\mathbf{z}_{t},t)\,.$ (8)
This step corresponds to estimating $\nabla\log p(\mathbf{r}_{t})$ within our conceptual velocity loss decomposition (where we separate representation learning from generation, as detailed in  Appendix A), sometimes termed the Representation Score Song et al. (2020). Intuitively, the model should discern salient patterns (e.g., object shapes, style characteristics, or conditioning signals) before denoising. Under the sufficient statistic assumption (see Assump. 1), the representation $\mathbf{r}_{t}$ effectively captures the essential information from the latent $\mathbf{z}_{t}$, allowing us to formulate the representation learning process as an independent objective:
$\mathcal{L}_{\mathrm{L2R}}\;=\;\mathbb{E}_{t,\mathbf{z}_{t}}\bigl{[}\|\,\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}(\mathbf{z}_{t},t)\!-\!\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})\,\|^{2}\bigr{]}\,,$ (9)
@@
In the final part of each reverse diffusion update in (3), known as the R2G stage, the extracted semantic representation is transformed into an updated latent with reduced noise:
@@
$\mathbf{r}_{t}=\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}(\mathbf{z}_{t},t)\quad\longrightarrow\quad\mathbf{z}_{t-\Delta t}=\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}(\mathbf{r}_{t},t)$ (12)
Hence, every time step naturally splits into (i) L2R for refining the representation and (ii) R2G for synthesizing an updated latent. This loop neatly implements the principle of “first representation, then generation”. Empirically, prior work (Yu et al., 2024; Xiang et al., 2023) confirms that early layers of the diffusion model predominantly focus on representation extraction, whereas later layers emphasize generative refinement. Consequently, the staged design mirrors the reverse-time diffusion trajectory, concluding in a final latent $\mathbf{z}_{0}$ that is decoded via $\mathcal{D}_{\bm{\theta}}$ to yield the synthesized output $\mathbf{x}_{0}$.